OpenStudy (anonymous):

Quadratic Equation: y=2x squared - 4x-1

OpenStudy (anonymous):

2x^2 -4x -1 x1,2 = (4 +/- (16 + 8)^0.5) / 4 x1,2 = 1 +/- ((6)^0.5)/2

jhonyy9 (jhonyy9):

\[y=2x ^{2}-4x-1\]

jhonyy9 (jhonyy9):

x_1,_2=(4+/-sqrt(16+8))/4=(4+/-2sqrt(6))/4=(2+/-sqrt(6))/2